Answer this question please with solution Thank you.




Two point charges, q1 = 8 µC and q2 = -5 µC, are separated by a distance r = 0.1 m. What is the magnitude of the electric force? Note: 1 µC = 10-6 C.

1
Expert's answer
2022-03-09T10:32:59-0500

The Coulomb's law says


"F=k\frac{q_1q_2}{r^2}""F_1=9*10^9*\frac{8*10^{-6}*5*10^{-6}}{(0.1)^2}=36\:\rm N"